OpenStudy (anonymous):

Derivative of sqrt(7x-8)

OpenStudy (stamp):

\[f(x)=\sqrt{7x-8}\]\[f'(x)=\frac{1}{2\sqrt{7x-8}}d(7x-8)\]\[=\frac{7}{2\sqrt{7x-8}}\]Ch-ch-chainrule

OpenStudy (anonymous):

Thank you so much